Find the amplitude, phase shift, and period of the function.
y = 1-2sin(2x + pi/3)
Give the exact values, not decimal approximations.

Answer by stanbon(75887)

Find the amplitude, phase shift, and period
of the function.
y = 1-2sin(2x + pi/3)
Give the exact values, not decimal approximations.
-----
Form:: y = a*sin(b(x-c)) + d
amp = |a|
period = (2pi)/b
phase shift = c/a
-------------------------
Your Problem::
y = -sin(2(x-(-pi/6)) + 1
a = -1 ; b = 2 ; c = -pi/6 ; d = 1
---
amp = |-1| = 1
period = (2pi)/2 = pi
phase shift = c/a = pi/6
==========
Cheers,
Stan H.
-----------